OpenStudy (anonymous):

improper integral from 0 to infinity (dtheta/(1+e^theta))

OpenStudy (anonymous):

hint\[\frac{1}{1+e^{\theta}}=\frac{1+e^{\theta}-e^{\theta}}{1+e^{\theta}}=1-\frac{e^{\theta}}{1+e^{\theta}}\]